OpenStudy (anonymous):

Solve sin(x)=cos(pi/5), where 3pi2

OpenStudy (anonymous):

is that 3pi/2??

OpenStudy (anonymous):

yes and its sin(x)=-cos(pi/5) forgot - infront of cos

OpenStudy (anonymous):

okay.. cos pi/5 = -sin (3pi/2+ pi/5) add 3pi/5 and pi/2, that is 17pi/10

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

sinx=-sin(17pi/10)

OpenStudy (raden):

sin(x) = -cos(pi/5) sin(x) = -sin(pi/2 - pi/5) sin(x) = sin(-(pi/2 - pi/5)) sin(x) = sin(pi/5 - pi/2) use the identity : sin(x) = sin(k*2pi + x) so, the equation above can be sin(x) = sin(k*2pi + pi/5 - pi/2) cancel out the sin from both sides x = sin(k*2pi + pi/5 - pi/2) for k = 0, we get x = pi/5 - pi/2 (not satisfies) for k = 1, we get x = 2pi + pi/5 - pi/2 =17pi/10 it is satisfies for x in the 3rd quadrant

OpenStudy (raden):

i mean in the 4th quadrant :)

OpenStudy (anonymous):

sin(x) = -sin(pi/2 - pi/5) sin(x) = sin(-(pi/2 - pi/5)) instead of -sin(x)=sin(-(x)) can you use sin(pi+(x))?

OpenStudy (raden):

i used the identity of sinx = cos(pi/2 - x)

OpenStudy (anonymous):

no for the line -sin -> sin

OpenStudy (raden):

oh, i used identity -sin(x) = sin(-x)

OpenStudy (anonymous):

yes can u use -sin(x)=sin(pi+x) instead

OpenStudy (raden):

yes, that will works too

OpenStudy (anonymous):

how about when you used sin(x) = sin(k*2pi + x) , sin(x)=sin(pi-x+k^2pi) works too right

OpenStudy (raden):

it's thought one

OpenStudy (anonymous):

?? what thought one